Two samples of a monatomic ideal gas are in separate containers at the same conditions of pressure, volume, and temperature ( please help!!!
Answer #1

Sample 1: (P1,V1) \rightarrow (P1, V2) \rightarrow (P2, V2)

Work done (w) = w1 + w2

= - Pext(Vf + Vi) + 0 = - 1.00atm(2.00 - 1.00)L = -1.00atm-L

Sample 2: (P1, V1) \rightarrow (P2, V1) \rightarrow (P2, V2)

Work done = w1 + w2 = 0 + (-Pext(V2 - V1))

= - 2.00atm(2.00 - 1.00)L = - 2.00atm-L

Difference of work between first and second sample = -1.00atm-L - ( -2.00atm-L)

= 1.00atm-L. (Answer)